OpenStudy (anonymous):

What is the first step in solving the equation 4x + 2y = 5 for x? A. Subtract 2y from both sides of the equation. B. Divide both sides of the equation by 5. C. Divide both sides of the equation by 4. D. Subtract 5 from both sides of the equation.

OpenStudy (anonymous):

To solve the equation, what do you need to do?

OpenStudy (anonymous):

solve for x and y

OpenStudy (anonymous):

But you want so solve it for x only (in this case), to obtain the x value, you must have an expression like x=~~ So, what would you do first

OpenStudy (anonymous):

ill give medals and become fan for first right response

OpenStudy (math&ing001):

@courtney-ray Learning how to do it is more important than getting the right response from someone. To solve for 'x' in any equation, we isolate x on one side and the other terms in the other side like this: 4x + 2y = 5 - 2y = - 2y 4x = 5 - 2y 1/4 * (4x = 5 - 2y) x = 5/4 - 2/4 y x = 5/4 - (1/2)y

OpenStudy (anonymous):

thank you